# Stereo Hand-Object Reconstruction for Human-to-Robot Handover

### v1

@@ Abstract
Jointly estimating hand and object shape ensures the success of the robot grasp in human-to-robot handovers. However, relying on hand-crafted prior knowledge about the geometric structure of the object fails when generalising to unseen objects, and depth sensors fail to detect transparent objects such as drinking glasses. In this work, we propose a stereo-based method for hand-object reconstruction that combines single-view reconstructions probabilistically to form a coherent stereo reconstruction. We learn 3D shape priors from a large synthetic hand-object dataset to ensure that our method is generalisable, and use RGB inputs instead of depth as RGB can better capture transparent objects. We show that our method achieves a lower object Chamfer distance compared to existing RGB based hand-object reconstruction methods on single view and stereo settings. We process the reconstructed hand-object shape with a projection-based outlier removal step and use the output to guide a human-to-robot handover pipeline with wide-baseline stereo RGB cameras. Our hand-object reconstruction enables a robot to successfully receive a diverse range of household objects from the human.
@@ I Introduction
As humans and robots work together more closely [1], exchanging objects is one of the essential skills to enable smooth collaboration between humans and robots. In particular, human-to-robot handover is the exchange of objects between a human giver and a robot receiver [2]. In this highly interactive scenario, hand-object shape reconstruction is an essential task as the object shape informs the robot about how to grasp the object, while the hand shape allows the robot to avoid potentially dangerous contact with the human.
Visual modalities, including RGB and depth, are used in hand-object shape reconstruction for human-to-robot handover. Previous works [3, 4, 5, 6, 7] used depth sensors to capture the pointcloud of the scene and applied segmentation models to extract the pointcloud of the hand and object. Depth sensors can capture opaque objects but not transparent objects in the scene [8], causing the robot to fail to handover objects such as glasses. To accommodate transparent objects, some works have attempted to perform human-to-robot handovers with wide-baseline stereo RGB cameras [9]. However, these works can only handle container-like objects as they assume that the object is upright and rotationally symmetric.
@@
We consider the scenario where a person hands over various objects to a robot and the robot aims to grasp the object successfully while avoiding the human, and place the object at a predefined delivery location (Fig. 1). Container-like objects, such as drinking cups and glasses, can be empty or filled with content. The presence of content can alter the appearance of transparent containers which affects the performance of depth-based methods. Unlike depth cameras, RGB cameras can capture the appearance of generic objects, including transparent objects. Moreover, capturing stereo views instead of a single view can help tackle the issue of occlusion by increasing the visible area of the hand and object to the model [10].
We propose StereoHO [^1] Code and videos of the experiments available at https://qm-ipalab.github.io/StereoHO/, a hand-object reconstruction model for wide-baseline stereo inputs. Our method quantifies the uncertainty of single-view predictions, and take the uncertainty into account when forming the stereo reconstruction. For each local 3D patch, we predict the shape as a probability distribution over a set of learned discrete shape embeddings. A high variance represents an uncertain prediction that may be caused by issues such as occlusion. Predictions from the left and right views are then combined via element-wise multiplication and decoded jointly to form the stereo reconstruction. We train our model entirely on synthetic 3D hand-object data [11] with a large amount of objects and without specific assumptions about the object’s geometry to ensure its generalisation ability.
We then present the first human-to-robot handover method that jointly reconstructs hand and object shape from stereo RGB inputs. We apply our proposed reconstruction method on wide-baseline stereo cameras and track the hand’s movement over time. To ensure multi-view consistency, the reconstructed pointcloud is projected back to the left and right camera views and points outside of the hand and object segmentation masks are removed.
Experiments show that our hand-object reconstruction method outperforms previous methods in single-view and stereo settings for object reconstruction, while remaining competitive for hand reconstruction in terms of Chamfer distance. When integrated into a human-to-robot handover pipeline, our method can successfully handover objects of diverse shapes and appearances, including filled/empty transparent containers, boxes and thin objects like CDs.
@@ II Related work
II-A Perception for human-to-robot handover
Recent works in human-to-robot handover  [4, 5, 6, 7, 9, 12, 3] focused on markerless settings as it is more applicable to in-the-wild human-to-robot interactions. Human grasps were classified into different types based on a human grasp taxonomy to select the corresponding predefined safe robot grasp direction and enable the robot to receive the object [3]. This approach limits the grasp types that the robot can handle as the high degree of hand articulation implies that there is a continuous distribution of many possible ways for a human to grasp an object. Moreover, to handle a diverse set of objects, the object shape should be considered.
@@ II-A Perception for human-to-robot handover
Later works improved the perception by capturing the shape of the hand and the object [4, 5, 6, 7, 9, 12], and used existing grasp estimation methods to predict possible grasps on the object shape. Grasps that would cause a collision between the robot and the human are removed for safety. A two-stage system was proposed to perform human-to-robot handovers of everyday objects [4]. The human and the object are first segmented from the input RGB image. The depth image is then lifted into a pointcloud and the background is removed using the segmentation masks. Finally, an imaginary plane is added in the grasping direction to approximate the task as a top-down grasp estimation problem using GG-CNN [13]. To enable more natural handovers [5], 6-DoF grasps are used to perform the handover to allow more freedom in the robot’s movement.
Pointcloud is a common representation for the shape of the human hand and the object in existing methods for human-to-robot handovers [4, 5, 6, 7], as pointclouds can be directly obtained from a depth sensor and grasp estimation methods rely on this representation [13, 14]. Although pointclouds retrieved from a depth sensor is accurate enough for human-to-robot handover, the shape retrieved is dependent on the view direction. The object shape is often incomplete due to occlusions, affecting the grasping performance. Depth sensors cannot reliably detect transparent and reflective objects, causing failure in the grasp estimation. Depth restoration overcomes this issue and completes the pointcloud of transparent objects [15]. However, a depth camera is still required which is not as commonly available as RGB cameras.
Figure 2: Depth-based sensing fails on transparent objects (left), while relying on shape priors is not generalisable (right).
@@
Other methods use stereo RGB images as input to reconstruct the shape of the object [9, 12]. The shape of containers, such as drinking glasses and cups, was reconstructed from stereo RGB images by assuming a rotationally symmetric geometry and exploiting the predicted segmentation masks of the object [16]. This enables human-to-robot handover by grasping at the centroid of the object [9]. Safety of the human can be considered by estimating the 3D hand keypoints [12]. Robot grasps are sampled along the vertical axis of the object and grasp points that are too close to the hand keypoints are removed. However, these methods do not estimate 6-DoF grasps and thus limit the naturalness of the robot movement as the object must maintain an upright orientation during the handover. The assumption on the object shape can also be easily invalidated by non-container objects, causing failure in the object shape estimation. Furthermore, estimating the shape of both the hand and the object provides more information for the human’s safety.
II-B Hand-object reconstruction from RGB images
@@
Joint hand-object reconstruction from RGB images in an interaction scenario is challenging due to severe mutual occlusions between the hand and the object [11]. Large hand-object interaction datasets, both synthetic and collected in real environments [11, 17, 18], have helped improve the reconstruction quality. Existing methods can be grouped into single-view and multi-view, and multi-view methods can be further split into dense and sparse. Single-view reconstruction methods [11, 19, 20, 21] learn 3D shape priors from hand-object datasets, but remains challenging as the method has to complete the shape of the unobserved parts of the object. Multi-view methods [22, 23, 24, 25, 10] overcome the problems of occlusion and unobserved object parts by increasing the number of input views. In particular, dense multi-view methods [22, 23, 24, 25] can produce high quality reconstructions but require a large number of input views and computationally expensive optimisation strategies, limiting the applicability in reactive scenarios such as human-to-robot handovers. Sparse multi-view methods [10] balances the reconstruction quality of dense methods with the speed of single-view methods. However, training only on synthetic data can affect the model’s ability to generalise to real world scenarios (sim-to-real gap) [10].
@@
Figure 3: Our proposed hand-object reconstruction method with two cropped images from a wide-baseline stereo camera. StereoHO first performs shape estimation from each view independently to obtain the predicted probability distributions, $P_{L}=\{P_{HL},P_{OL}\}$ and $P_{R}=\{P_{HR},P_{OR}\}$, over the shape codebooks. The stereo-based probability distribution $P=\{P_{H},P_{O}\}$ is computed by element-wise multiplication of $P_{L}$ and $P_{R}$. The trained SDF decoder transforms the stereo prediction into the hand-object TSDF. Surface points, sampled as a pointcloud $\mathcal{P}$ from the TSDF, are projected into each view using the predicted camera projection parameters. We use the segmentation masks to remove the outliers to obtain the final pointcloud $\mathcal{P}^{\prime}$. KEYS – $K$: intrinsics calibration parameters, $L$: left view, $R$: right view; $H$: hand, $O$: object.
@@ III Proposed method
In this section, we introduce StereoHO for stereo hand-object reconstruction. We detail how we tackle the sim-to-real gap when training on synthetic data, highlight how our method combines single-view predictions for stereo reconstruction, and present our human-to-robot handover pipeline.
@@ III-A Stereo-based RGB hand-object reconstruction
Given stereo RGB images $I$ cropped and centred on the hand, wrist poses $T_{H}$ and hand-object segmentation masks $M\in\{M_{H},M_{O}\}$ in each view, our method predicts the pointcloud $\mathcal{P}\in\{\mathcal{P}_{H},\mathcal{P}_{O}\}$ representing the surface of the hand and object jointly over both views. We task the network to predict the signed distance fields (SDF) of the hand and object as an intermediate step. Compared to pointclouds, SDF represents smooth continuous surfaces and is easier to learn [26]. We train our model only on synthetic data. As segmentation masks are domain-invariant, and to facilitate sim-to-real transfer, we include hand and object segmentation masks as input to the model. The object segmentation mask allows our model to avoid erroneously including background clutter in the reconstructed shape.
To combine individual predictions for a coherent stereo reconstruction, the model should consider and weigh their uncertainty. For example, uncertainty should be higher for areas that are occluded and lower for areas that are clearly visible. To achieve this, we treat the hand-object shape reconstruction as a local shape classification task based on a learned discrete shape codebook. The output is a predicted probability distribution $P$ over the discrete codes where the variance of the distribution represents the uncertainty of the prediction. We learn this discrete shape codebook by training a 3D autoencoder with vector quantization to obtain quantized shape embeddings on a synthetic hand-object dataset [11]. We train an image encoder to predict the probability distribution of shape codes for each local 3D patch from the individual input views and the learned shape codebook. The predicted probability distributions from each view are combined into a single prediction via element-wise multiplication. We obtain the most likely shape code at each local 3D patch and decode the global shape jointly to ensure a coherent hand-object shape.
@@ III-B Learning discrete 3D shape embeddings
To learn the 3D shape embeddings, we train a 3D P-VQ-VAE autoencoder [27] to encode the hand and object truncated signed distance field (TSDF) independently. We extract the TSDF of each training sample and use 3D convolution layers to encode the shape of local patches. We also use the vector quantization process to obtain a codebook of 3D shape embeddings. The SDF decoder applies 3D convolution layers to the 3D shape embeddings jointly to retrieve the SDF values. We train the autoencoder using the following loss function,
$\mathcal{L}_{ae}=||s-\hat{s}||+||sg[z_{e}]-e||^{2}_{2}+\beta||z_{e}-sg[e]||^{2}_{2},$ (1)
where $s$ and $\hat{s}$ are the ground truth and predicted SDF values, $e$ is the vector quantized shape embedding, $z_{e}$ is the shape embedding output by the encoder before vector quantization, $sg$ is the stop gradient operator, and $\beta$ is the hyper-parameter controlling the scale of the commitment loss.
Figure 4: Our proposed robot control pipeline for human-to-robot handover (modules from other works are in grey). We first perform hand-object detection to obtain bounding boxes $B$. Hand-object segmentation masks $M$ and wrist poses $T_{H}$ are estimated on the image cropped around the hand. We combine the outputs from the preprocessing steps for stereo hand-object reconstruction to obtain pointcloud $\mathcal{P}^{\prime}$. Grasp estimation is performed on the reconstructed shape and transformed at each timestep using the wrist poses $T_{HB}$. The wrist pose in robot coordinate space $T_{HB}=T_{WB}*T_{HW}$ is calculated using the world to robot base transform $T_{WB}$ obtained using the hand-eye calibration process [28].
III-C From single-view predictions to stereo reconstruction
We reconstruct the hand and object shape from each view by predicting the learned 3D shape embeddings for each local patch. We use the trained encoder to extract the 3D shape embeddings for each training sample and the index corresponding to the codebook entry. We treat shape estimation as a classification task by predicting the codebook index corresponding to the shape embedding at each local 3D patch. We use the following weighted cross-entropy loss,
$\mathcal{L}_{ce}(p,p_{gt})=-\sum_{c}^{C}{w_{c}}\cdot p_{gt}(c)\cdot\text{log }p(c),$ (2)
where $p$ and $p_{gt}$ are the predicted and ground truth probability for index $c$, $C$ is the codebook size and $w_{c}$ is the weight for the index $c$.
We build a canonical 3D representation by constructing a 3D grid centred on the hand wrist and retrieve the corresponding image embedding for each local patch by projecting the 3D position to 2D image space using the wrist pose $T_{HL}$ and $T_{HR}$. The resulting 3D representation is processed with 3D convolution layers and the probability distribution $P_{L}$ and $P_{R}$ over codebook size $C$ is predicted per local patch for the left and right views, respectively.
To obtain the overall stereo prediction, we multiply the predicted probability distributions from the left and right views: $P=P_{L}\cdot P_{R}$. To obtain the pointcloud of the hand and object $\mathcal{P}_{H}$ and $\mathcal{P}_{O}$, we extract TSDF values in a grid using the trained SDF decoder and keep the points within a threshold SDF value of 1 cm to the surface.

### v2

@@ Abstract
Jointly estimating hand and object shape facilitates the grasping task in human-to-robot handovers. However, relying on hand-crafted prior knowledge about the geometric structure of the object fails when generalising to unseen objects, and depth sensors fail to detect transparent objects such as drinking glasses. In this work, we propose a stereo-based method for hand-object reconstruction that combines single-view reconstructions probabilistically to form a coherent stereo reconstruction. We learn 3D shape priors from a large synthetic hand-object dataset to ensure that our method is generalisable, and use RGB inputs to better capture transparent objects. We show that our method reduces the object Chamfer distance compared to existing RGB based hand-object reconstruction methods on single view and stereo settings. We process the reconstructed hand-object shape with a projection-based outlier removal step and use the output to guide a human-to-robot handover pipeline with wide-baseline stereo RGB cameras. Our hand-object reconstruction enables a robot to successfully receive a diverse range of household objects from the human.
@@ I Introduction
As humans and robots share the same workspace [1], exchanging objects is one of the essential skills to enable smooth collaboration. We consider the scenario where a human hands over various everyday objects to a robot and the robot aims to grasp the object while avoiding the human (human-to-robot handover [2, 3, 4]), and place the object at a predefined delivery location (Fig. 1). Objects can vary in their physical properties, such as appearance, transparency, shape, size, and mass [5]. Container-like objects, such as cups, drinking glasses, or boxes, can be empty or filled with content, affecting the perception of the appearance or the weight of the object [4, 6]. In this highly interactive scenario, hand-object shape reconstruction is an essential task as the object shape informs the robot about how to grasp the object, while the hand shape allows the robot to avoid dangerous contact with the human.
Visual modalities (RGB and depth images) are used in hand-object shape reconstruction for human-to-robot handover. Previous works [7, 3, 8, 9, 10] used depth sensors to capture the pointcloud of the scene and applied segmentation models to extract the pointcloud of the hand and object. However, transparent objects are not easy to be captured by depth sensors [11] and object shapes are incomplete due to partial visibility or occlusions in the scene when using a single camera [3]. Therefore, depth completion methods are devised to fill in the missing depth values [12], but the pointcloud is only filled in on the visible side of the object and a depth camera is still required. As RGB-only cameras are more commonly available than RGB-D cameras, and less expensive, some works use wide-baseline stereo RGB cameras to perform human-to-robot handovers of container-like objects and address partial visibility and occlusions of the object [4, 6]. However, these works are limited to only container-like objects (rotationally symmetric shape), also assuming the object is held upright during the handover due to the presence of a content.
@@
We propose StereoHO, a hand-object reconstruction method for wide-baseline stereo inputs. Our method quantifies the uncertainty of single-view predictions, e.g. caused by self-occlusions, by predicting low-dimensional embeddings of hand and object shapes as a probability distribution over learned codebooks of discrete shape embeddings. Individual predictions from each camera views are aggregated and decoded jointly to form a coherent stereo reconstruction. We train our model on synthetic 3D hand-object data [13] with a large amount of objects and without specific assumptions about the object’s geometry to ensure its generalisation ability. We integrate StereoHO in the first human-to-robot handover pipeline that jointly reconstructs hand and object shapes from stereo RGB inputs while tracking the hand over time. The pipeline also ensures multi-view consistency and reconstruction quality for accurate robot grasp predictions. Experiments [^1] Code and videos of the experiments available at https://qm-ipalab.github.io/StereoHO/ show that StereoHO outperforms previous methods in single-view and stereo settings for object reconstruction, while remaining competitive for hand reconstruction in terms of Chamfer distance. Furthermore, our pipeline can successfully handover objects of diverse shapes and appearances, including filled/empty transparent containers, boxes and thin objects.
@@ II Related work
Recent works in human-to-robot handover  [3, 8, 9, 10, 4, 6, 7] focused on markerless settings that are more applicable to daily human-to-robot interactions. Human grasps were classified into different types to select the corresponding predefined safe robot grasp direction and enable the robot to receive the object [7]. However, this approach does not consider diverse object shapes and limits the grasp types that the robot can handle. Later works improved the perception by capturing the shape of the hand and the object [3, 8, 9, 10, 4, 6], and used existing grasp estimation methods to predict possible grasps on the object shape. Grasps that would cause a collision between the robot and the human are removed for safety. A two-stage system was proposed to perform human-to-robot handovers of everyday objects [3]. The first stage segments the person and the object from the input RGB image. The depth image is lifted into a pointcloud and the background is removed using the segmentation masks. The second stage adds an imaginary plane in the grasping direction to approximate the task as a top-down grasp estimation problem using GG-CNN [14]. Using 6-DoF grasps allows more freedom in the robot’s movement, enabling more natural handovers [8].
Existing methods for human-to-robot handovers [3, 8, 9, 10] and grasp estimation [14, 15], rely on pointcloud as a representation for hand and object shape. However, pointclouds of objects are view-dependent and are incomplete in the presence of occlusions [3] or when the objects are transparent and reflective [11]. Incomplete object shapes or missing pointclouds for transparent objects (Fig. 2) can results in inaccurate grasp estimations. To overcome this issue, depth restoration is a method to complete the pointcloud of transparent objects [11]. However, a depth camera is required which is not as commonly available as RGB cameras. Moreover, the completed pointcloud often only consists of the visible side of the object [12]. An alternative representation for object shapes is Signed Distance Dields (SDFs) that encode into a vector the signed distance between a set of points in 3D and their closest point on the 3D object [16]. As an implicit representation of smooth and continuous surfaces, SDFs distinguish the inside and outside parts of an arbitrary object and are easier to learn compared to pointclouds [16]. None of the existing works has used SDFs for reconstructing object-only or hand-object shapes for more accurate human-to-robot handovers.
@@
Other methods use stereo RGB images as input to reconstruct the shape of the object [4, 6]. The shape of containers, such as drinking glasses and cups, was reconstructed from stereo RGB images by assuming a rotationally symmetric geometry (Fig. 2) and exploiting the predicted segmentation masks of the object [17]. This enables human-to-robot handover by grasping at the centroid of the object [4]. Human safety is considered by estimating the 3D hand keypoints [6]. Robot grasps are sampled along the vertical axis of the object and grasp points that are too close to the hand keypoints are removed. However, these methods do not estimate 6-DoF grasps and thus limit the naturalness of the robot movement as the object must maintain an upright orientation during the handover. The assumption on the object shape is easily invalidated by non-container objects, causing failure in the object shape estimation.
Figure 2: Limitations of existing human-to-robot handover approaches. Depth-based sensing fails on transparent objects (left). Relying on shape priors does not generalise (right).
@@
Joint hand-object reconstruction from RGB images in an interaction scenario is challenging due to severe mutual occlusions between the hand and the object [13]. Large hand-object interaction datasets, both synthetic and collected in real environments [13, 18, 19], have helped improve the reconstruction quality. Existing methods can be grouped into single-view and multi-view, and multi-view methods can be further split into dense and sparse. Single-view reconstruction methods [13, 20, 21, 22] learn 3D shape priors from hand-object datasets, but completing the shape of the unobserved parts of an object remains challenging. Multi-view methods [23, 24, 25, 26, 27] overcome the problems of occlusion and unobserved object parts by increasing the number of input views. Dense multi-view methods [23, 24, 25, 26] use a large number of input views to produce high quality reconstructions but are computationally expensive, limiting the applicability in reactive scenarios. Sparse multi-view methods [27] balances the reconstruction quality of dense methods with the speed of single-view methods. However, training only on synthetic data can affect the model’s ability to generalise to real world scenarios (sim-to-real gap) [27].
@@
Figure 3: Our proposed method, StereoHO, for hand-object reconstruction with two cropped images from a wide-baseline stereo camera. StereoHO predicts the probability distributions, $P_{L}=[P_{HL},P_{OL}]$ and $P_{R}=[P_{HR},P_{OR}]$, over the shape codebooks for each view and combined them into a coherent probability distribution $P=[P_{H},P_{O}]$. The trained SDF decoder transforms $P$ into the hand-object T-SDF. Sampled surface points from the T-SDF (pointcloud $\mathcal{P}$) are projected into each view using the predicted camera projection parameters and outliers are removed by using the segmentation masks to obtain the final pointcloud $\mathcal{P}^{\prime}$. KEYS – $K$: intrinsics calibration parameters, $L$: left view, $R$: right view; $H$: hand, $O$: object.
@@ III Proposed method
In this section, we introduce StereoHO for stereo hand-object reconstruction and our human-to-robot handover pipeline based on StereoHO.
@@ III-A Stereo-based RGB hand-object reconstruction
We design StereoHO to predict the joint surface of the hand ($H$) and object ($O$), represented as pointcloud $\mathcal{P}=[\mathcal{P}_{H},\mathcal{P}_{O}]$, from the RGB images of a stereo-based camera. Images are cropped and centred on the person’s hand holding an object. We formulate stereo-based hand-object shape reconstruction as a view-dependant image-to-shape classification task (encoding) followed by an aggregation step [20] and a transformation of the aggregated shape codes into the final joint hand-object pointcloud (decoding). Fig. 3 illustrates the inference pipeline of StereoHO. To avoid erroneously including background clutter in the reconstructed shape, we include the segmentation masks of the object, $M_{O}$, and the hand, $M_{H}$, as input to the model. We also predict the wrist pose $T_{H}$ for each view to relate the 3D shapes with the cameras views, assuming the camera poses are not available.
For the encoding part, we define a 3D grid centred on the hand wrist $T_{Hv}$, where $v\in\{L,R\}$ is the camera view. For each view $v$, we project the 3D position of each voxel into the image space using the wrist pose and associate each voxel with an image embedding, extracted with a convolutional neural network (e.g., ResNet-18 [28]), at the corresponding pixel location [20]. We use 3D convolutional layers to transform the resulting 4D tensor into a probability distribution $P_{v}\in[0,1]^{D\times D\times D\times C}$ with respect to a codebook of $C$ shape embeddings, $\mathcal{C}=\{e_{c}\in\mathbb{R}^{S}\}_{c=1}^{C}$. We combine the predicted probability distributions from each view into a single coherent prediction via element-wise multiplication: $P=P_{L}P_{R}$. We thus define another 4D tensor of shape embeddings that are selected from $\mathcal{C}$ based on the index of the highest probability for each element of the grid, $E\in\mathbb{R}^{D\times D\times D\times S}$. The same procedure is performed independently for the hand and the object, resulting in the two shape-based tensors, $E_{O}$ and $E_{H}$, respectively.
For the decoding part, we apply hand and object specialised networks, consisting of 3D upconvolutional layers, to obtain the pointclouds $\mathcal{P}_{H}$ and $\mathcal{P}_{O}$ from $E_{O}$ and $E_{H}$, respectively. Each network (SDF decoder) predict SDF values, $\hat{s}\in\mathbb{R}^{D\times D\times D}$. We then sample the Truncated-SDF (T-SDF) at discrete positions and only the values within a threshold of 1 cm to the surface are kept as the predicted pointclouds. We ensure multi-view consistency by re-projecting the predicted hand and object pointcloud $\mathcal{P}_{H}$ and $\mathcal{P}_{O}$ in the images of both camera views using the wrist poses $T_{Hv}$, and by filtering out any outlier $\mathcal{P}_{B}$ with respect to the segmentation masks $M_{H}$ and $M_{O}$, resulting in the final prediction $\mathcal{P}^{\prime}$.
Next, we detail how we learn the codebooks of 3D shape embeddings and the neural network to reconstruct the hand-object pointcloud from the shape embeddings, and the image-to-shape embeddings. As hand-object data is hard to obtain and annotate, we train the components of StereoHO only on synthetic data. Note that the use of segmentation masks facilitate sim-to-real transfer as they are domain-invariant.
Figure 4: Our proposed pipeline for human-to-robot handover. For each frame, hand-object detection extracts bounding boxes $B$ to crop the input images around the hand and then used to segment hand-object masks $M$ and estimate wrist poses $T_{H}$. Our StereoHO uses these outputs along with the image crops to reconstruct the hand-object shape $\mathcal{P}^{\prime}$. We estimate the grasp $g$ on the reconstructed shape and transform it into the robot coordinate system using the wrist poses $T_{HB}$.
@@ III-B Learning discrete 3D shape embeddings
We learn two discrete codebooks, one for the objects $\mathcal{C}_{O}$ and one for the hands $\mathcal{C}_{H}$, by training a 3D Patch-wise Encoding Variational Autoencoder with Vector Quantization [29] on a synthetic hand-object dataset [13]. For each shape (object or hand) in the training set, we transform the shape representation from a 3D mesh to SDFs, discretized in a 3D grid (T-SDFs), $s\in\mathbb{R}^{D\times D\times D}$. The model encodes local patches of the T-SDF with 3D convolutional layers into a shape code $z_{e}\in\mathbb{R}^{S}$, maps each shape code to the closest quantized shape embedding $e\in\mathbb{R}^{S}$ in the codebook (vector quantization [30]), and reconstructs the shape $\hat{s}$. Through mapping the continuous embeddings into a finite set of learned embeddings, vector quantization reduces the high-dimensionality of the shape embeddings [30]. The SDF decoder then retrieves the T-SDF values $\hat{s}$ from $e$.
We use the following loss to train each model by simultaneously learning the autoencoder [^2] We tested the two autoencoders on ObMan [13], achieving a Chamfer distance of 0.015 cm2 and 0.437 cm2 for the hand and object, respectively. parameters and codebook,
$\mathcal{L}_{ae}=|s-\hat{s}|+||sg[z_{e}]-e||^{2}_{2}+\beta||z_{e}-sg[e]||^{2}_{2},$ (1)
where $sg$ is the stop gradient operator. The hyper-parameter $\beta$ controls the scale of the commitment loss (last term), that helps stabilize the training [29]. We initialise the codebook with a uniform distribution. After training the models for object and hand, we only keep the trained SDF decoders and codebooks and link them to the image-to-shape encoder.
III-C Training of the image-to-shape encoder
We train an image-to-shape encoder that transforms the input image and segmentation masks into a probability distributions over the hand and object codebook. We use ResNet-18 [28] pre-trained on ImageNet[31] as our image encoder and 3D convolutional layers as prediction heads for the hand and object. We train all layers with a weighted cross-entropy loss function defined as
$\mathcal{L}_{ce}=-\sum_{c}^{C}{w_{c}}\cdot p_{gt}(c)\cdot\text{log }p(c),$ (2)
where $p_{gt}$ and $w_{c}$ are the ground-truth probability and weight for index $c$. We set a lower weight for the index $c$ corresponding to empty space as the majority of the space surrounding the hand and object in the 3D grid is empty.